Show that , for the DE
d²y/dx² + a(x) dy/dx+b(x)y=0

exp(mx) is a particular integral if m²+am+b=0 . Hence find the value of m so that exp(mx) is a particular integral of the equation
(x-2) d²y/dx² - (4x-7)dy/dx+(4x-6)y=0
1
Expert's answer
2019-08-19T13:36:55-0400

Let's make a substitution "y=exp(mx)"

"y'=my"

"y''=m^2y"

Hence

"y''+a(x)y'+b(x)y=m^2y+a(x)my+b(x)y"

Therefore, if

"m^2+am+b=0"

"y''+a(x)y'+b(x)y=0"

Which means y=exp(mx) is particular integral of the equation.  

In the equation "(x-2)y''-(4x-7)y'+(4x-6)y=0"

"a=-\frac{4x-7}{x-2};\quad b=\frac{4x-6}{x-2}"

Let's solve the eqation

"m^2+am+b=0"

"m^2-\frac{4x-7}{x-2}m+\frac{4x-6}{x-2}=0"

"m=\frac{4x-7-\sqrt{(4x-7)^2-4(x-2)(4x-6)}}{2(x-2)}=\frac{4x-7-1}{2(x-2)}=2"

therefore "y=e^{2x}" is a particular integral of the equation